Gavin purchased three types of meat at the deli totaling 5 lb. He bought
1 1/2 lb of bologna and 2 1/4 lb of ham. How much of the other type of meat die Gavin purchase?

Answer by checkley71(8403)

1.5+2.25+x=5
3.75+x=5
x=5-3.75
x=1.25 amount of the third type of meat.